OpenStudy (mstv1112):

Solve the equation. Show all steps. y – 5 = –8

TheSmartOne (thesmartone):

Just ad 5 on both sides. y=-8+5

TheSmartOne (thesmartone):

@mstv1112 You should get y=-3

TheSmartOne (thesmartone):

Steps y-5=-8 +5 +8 y=-8+5 y=-3

OpenStudy (mstv1112):

Thank you very much, I suck at math.

TheSmartOne (thesmartone):

No problem.

OpenStudy (alprincenofl):

y=8-5 y=3 I am not sure

TheSmartOne (thesmartone):

There is a minus in front of the 8 and you have to add the 5.

OpenStudy (alprincenofl):

I understand that you add 5 on both sides. What I did is changing the sides of the numbers so it changes the negative.

TheSmartOne (thesmartone):

If you want -8+5= -1(8-5)=-1(3)=-3

OpenStudy (mstv1112):

How about this one? Solve for n. n – 40 = –35 would it just be -35+40

OpenStudy (alprincenofl):

@TheSmartOne , you are true

TheSmartOne (thesmartone):

Yes -35+40

TheSmartOne (thesmartone):

Which would be +5

OpenStudy (mstv1112):

Lauren needs at most $80.00 for gas. How many twenty dollar bills does she need?

TheSmartOne (thesmartone):

80/20 = 4

OpenStudy (mstv1112):

Wow aha dumb question for me to ask I read that wrong XD Thanks (facepalm)

TheSmartOne (thesmartone):

^^